SERENDIP (L Band Flat Feed) Receiver Log

as of Saturday, 16-Dec-2017 22:26:29 AST

Go to Most Recent Entry


NameEmailDateComment
ANS anolla@naic.edu 11/16/99 The 21 cm flatfeed antenna removed from receiver due to Hurricane Lenny.
ANS anolla@naic.edu 11/24/99 21 cm flatfeed receiver back to normal.The linefeed was disconected due to hurricane Lenny.Tsys. 64 deg.
ANS anolla@naic.edu 1/11/2000 On 1/3/2000 JR turned off the cryo system due that sys. warmed up cause of contamination.Receiver RF is on. Replace crossheads , and start pumping Dewar.
ANS anolla@naic.edu 1/12/2000 The crosshead ED 002 was replaced with crosshead ED 008. System cooled down. Temp. are 12.4 k deg at 16 stage and 54.6 k deg at 70 stage.
ANS anolla@naic.edu 2/24/2000 Dewar was intalled and cooled down.System was down since 2/7/2000 when Dewar warmed up due to small displacer was broken in two.Coco had a hard time taking it out of the cylinder.
ANS anolla@naic.edu 5/4/2000 The crosshead ED 002 was replaced with crosshead ED 026 .System was cooled down on 5/2/2000.Temp are 12.2 k deg at 16 deg stage and 54.7 k deg at 70 deg stage.
ANS anolla@naic.edu 8/29/2000 System cooled by J Rosa.
ANS anolla@naic.edu 9/8/2000 System not seen cal.Have to be checked. J Rosa found the cal cable broken.Need to be fixed.
ANS anolla@naic.edu 10/12/2001 System taken down for repairs , it has the bias box ac line shorted out due to water on it.
ANS anolla@naic.edu 1/23/2002 Working with a new system using a Miteq amp at room temp ( it has +/- 60 k deg at room temp.). LW taking measurements on it.
ANS anolla@naic.edu 2/1/2002 System installed yesterday. Was checked today for Tsys. Its cal is 4.35 ēk, the Tsys was +/- 85 ēk. Lisa double check it twice today and its Tsys was the same.
ANS anolla@naic.edu 3/4/2002 The recvr LO is causing RFI at 1160 Mhz , so we replaced the LO RG-9 coax cable from the rf cabinet to Hoffman box with a superflex one to see if the RG-9 shield was broken but instead the rfi signal increased by 5 db , so we suspect that is radiating thru the antenna .So a further test is need to pinpoint where the rfi is coming from. Also the LO box was taken down to check it's P.S. it looks to be intermittent.
ANS anolla@naic.edu 3/5/2002 LO chassis was installed back , it's P.S. was replaced with a new one.There was no time to check the recvr itself due that there was an experiment running.
LSWlwray@naic.eduThursday, November 7, 2002 17:05:36ANS noticed that the receiver was reading 280K tsys on Tuesday. We verified that it was still reading 280K this morning - it was. Tested the LNA, and it reads 80K instead of 40K. When the LNA was returned, the tsys went down to 115K. Tested system again in late afternoon - back to 280K. More testing tomorrow needed. System is up, but at reduced sensitivity.
ansanolla@naic.eduWednesday, November 13, 2002 08:56:33Receiver down in order to check an unstability on it.
LSWlwray@naic.eduWednesday, November 20, 2002 17:06:49Found LNA to be damaged (T=90K instead of 40K), receiver still down waiting for new Miteq amp. Possible cause of damage: lightening.
aansanolla@naic.eduFriday, December 20, 2002 12:47:21System back to service , its Tsys is +/- 100ēk. Cal value is 4ēk.The miteq amp was replaced by LSR.
ansanolla@naic.eduThursday, March 13, 2003 14:21:17There was a complaint that SETI recvr was not locking to our standard. We went up and found that the PLL at 1160 mhz was locked to our 5 mhz ref.Lisa is still checking it.
lswlwray@naic.eduThursday, March 13, 2003 14:31:24test
lswlwray@naic.eduThursday, March 13, 2003 15:23:37SETI receiver's upstairs LO (from Miteq PLO) is locked, verified with tests downstairs (radiated a birdie and checked). Also tested the second LO in the backend room. THat is OK too. SETI people noted that the receiver started working sometime around the time that we went upstairs to check the RF box. hmmm.... further test: radiated single birdie through Whip antenna at 1420.5 MHz for 12 minutes. SETI people see birdie OK. Wait until problem appears again. Possible temperature dependence of the PLO? Loose cable? Receiver functioning normally as of end of day.
ansanolla@naic.eduWednesday, March 19, 2003 13:32:36LSW found that recvr was out of lock. A bad cable between the 5 mhz buffer out (1160 Mhz PLL LO ref) and the patch panel inside the rf cabinet at CH 1 was intermittent. Replaced with another spare one.
LSWlwray@naic.eduFriday, May 2, 2003 15:53:39April 18 (Saturday) - receiver not functioning. April 21 (Monday) found the switchable AC power bar to be at fault, this was powering the whole front end. Power bar replaced, and fixed and returned on 23rd. System back to normal on afternoon of April 21
ansanolla@naic.eduTuesday, December 30, 2003 13:15:12Reciver was off , it was turned up by JRV.Since when , we dont know.
lswlwray@naic.eduThursday, March 25, 2004 15:22:431160 MHz LO signal seen in LBW data from last night.
ansanolla@naic.edsuTuesday, March 30, 2004 16:10:22LSW removed the recvr in order to seal it from generating L.O. rfi radiation.It's Hoffman box seal is completly corroded due to weather.
ansanolla@naic.eduWednesday, April 7, 2004 12:02:08System was taken down on 3/29/2004 in order to check it due to the 1160 mhz rfi produce by it.LSW check it and found poor isolation at the door , it was all corroded due to weather exposure. The door rfi gasket replaced , it solved the rfi problem.
ansanolla@naic.eduWednesday, August 4, 2004 11:40:26System was off at the control room.Was checked and it had a Tsys of a 182 K deg.
ansanolla@naic.eduThursday, December 30, 2004 09:25:29Recvr. Tsys was 159ēk deg.
ansanolla@naic.eduFriday, June 2, 2006 11:28:20The CH #1 SERENDIP recvr been removed.SERENDIP people started using the ALFA receiver for the same purpose.So they brougt a new system to monitor thru the ALFA 7 channels.


Back to Top